OpenStudy (anonymous):

how do you find the side of an area of 361

OpenStudy (hlambach):

side of what?

OpenStudy (jhannybean):

what is the area you are looking at?

OpenStudy (anonymous):

square

OpenStudy (hlambach):

the formula of a square is: a = s^2 S0 if you know the area 361 = s^2 Can you do it from there?

OpenStudy (anonymous):

s^2 could you explain that part?

OpenStudy (hlambach):

s^2 means s x s

OpenStudy (jhannybean):

s^2 means you're multiplying 2 of the same sides, and because it's a square, you're multiplying the sides with eachother. in order to find one side you would have to take the square root of 361.

OpenStudy (hlambach):

but if: 361 = s^2 What do you need to do to get the ^2 to the other side?